OpenStudy (anonymous):

Quick question on boundary conditions: One of the problems the Prof assigned had these boundary conditions (laplace equation in polar coord):\[v(r,\theta)=0\]\[v(r,\alpha)=0\]\[v(1,\theta)=f(\theta)\]v continuous I'm confused on what to do with the fourth "v continuous" condition.

OpenStudy (anonymous):

what is your equation?

OpenStudy (anonymous):

The laplace equation in polar coordinates is: \[v_{rr}+\frac{1}{r}v_r+\frac{1}{r^2}v_{\theta \theta}=0\]

OpenStudy (anonymous):

if u want to solve it with 'separation of variables' method i think u will come up with something like this: \( v(r,\theta )=(A r^k+Br^{-k}) (C \cos k \theta+ D \cos k \theta ) \) u see this answer have a critical point at r=0 but v is continuous so u must let B=0

OpenStudy (anonymous):

Not sure how he wants us to solve it, we started fourier transform today, but I happened to miss class >.< Could that possibly tie in here somwhere?

OpenStudy (anonymous):

nope... i think its not connected with fourier transform...